OpenStudy (anonymous):

- 1 times ( 1 ) ---- 8 --- 3 7 A. 5 -2 --- 7 B. - 3 ---- 8 C. - 1 ---- 21 D. 1 --- 3

OpenStudy (babyslapmafro):

\[-\frac{ 1 }{ 3 }\times 8\frac{ 1 }{ 7 }\]

OpenStudy (babyslapmafro):

is that the problem?

OpenStudy (anonymous):

yes

OpenStudy (babyslapmafro):

None of the answers equal the problem you have given

OpenStudy (anonymous):

then u must divid it

OpenStudy (babyslapmafro):

Oh, you are supposed to find out if you need to multiply or divide to get one of the answers given?

OpenStudy (anonymous):

yea i think so im not sure i couldnt figure it out.

OpenStudy (babyslapmafro):

well if you divide it you won't get any of the answers either...